# Grasp2Grasp: Vision-Based Dexterous Grasp Translation via Schrödinger Bridges

### v1

@@ Abstract
We propose a new approach to vision-based dexterous grasp translation, which aims to transfer grasp intent across robotic hands with differing morphologies. Given a visual observation of a source hand grasping an object, our goal is to synthesize a functionally equivalent grasp for a target hand without requiring paired demonstrations or hand-specific simulations. We frame this problem as a stochastic transport between grasp distributions using the Schrödinger Bridge formalism. Our method learns to map between source and target latent grasp spaces via score and flow matching, conditioned on visual observations. To guide this translation, we introduce physics-informed cost functions that encode alignment in base pose, contact maps, wrench space, and manipulability. Experiments across diverse hand-object pairs demonstrate our approach generates stable, physically grounded grasps with strong generalization. This work enables semantic grasp transfer for heterogeneous manipulators and bridges vision-based grasping with probabilistic generative modeling.
@@ 1 Introduction
Robotic grasping remains a central challenge in autonomous manipulation, especially in the context of dexterous, multi-fingered hands. Unlike parallel-jaw grippers or simple suction devices, dexterous hands offer significantly more control and flexibility but also bring high-dimensional configuration spaces [2, 15] and non-trivial contact dynamics [61]. These properties pose difficulties [49] in data collection, modeling, and learning robust grasp strategies. While recent works [42, 48, 85, 8, 93] in data-driven grasp generation have shown success in generating feasible grasps for specific hand-object pairs, they often rely on large-scale datasets [42, 81, 88, 29] curated through computationally expensive simulation or real-world trials. Moreover, such models typically operate in a hand-specific setting, which requires retraining or significant fine-tuning to adapt to new robotic hands [37] and poses a practical bottleneck in the rapidly evolving landscape of robotic hardware.
This lack of generalization across different robotic hands severely limits the applicability of learned manipulation policies in real-world systems [91]. In particular, with the emergence of increasingly diverse robotic platforms, it becomes infeasible to collect training data or train models from scratch for every possible hand. Instead, it is desirable to reuse grasp knowledge acquired from one hand to infer meaningful grasp strategies for another. However, the morphological differences between hands mean that naive transfer, such as copying joint angles or end-effector poses, might lead to physically invalid or unstable grasps [64, 54, 1]. This motivates the need for a framework that can translate the intent behind a demonstrated grasp from a source hand to a new target hand, taking into account both the underlying object geometry and the structural differences in hand kinematics.
@@ 2 Related Work
Grasp Transfer is closely related to teleoperation and imitation learning [3], where the objective is to map motions or actions from a source agent (e.g., a human or a robot) to a target robot. Teleoperation methods [72, 74, 95, 33, 21, 66] often aim for one-to-one replication of joint trajectories or end-effector poses, which is generally effective when the source and target share similar kinematics. However, such approaches struggle when transferring grasps between manipulators with significantly different morphologies, as direct replication can lead to infeasible or suboptimal grasps [64, 54, 1]. More recent work, such as RobotFingerPrint [37], addresses this challenge by mapping grippers into a shared Unified Gripper Coordinate Space (UGCS) based on spherical coordinates to facilitate grasp transfer. While promising, this approach requires simulation-based preprocessing and detailed gripper models, which impose a strong prior that limits flexibility. Moreover, the UGCS approach focuses on point correspondences instead of explicitly preserving functional grasp intent. In contrast, our formulation aims to translate grasps in a distributional sense grounded in physical properties. It requires no hand-specific simulation preprocessing, which enables broader applicability and more semantically meaningful grasp transfer.
@@
Optimal Transport (OT) and Schrödinger Bridges (SB) have recently emerged as powerful frameworks in the machine learning community for modeling distributional transformations. OT [69, 86, 65, 75] provides a principled way to compute mappings between probability distributions that minimize a cost function, typically grounded in geometric distances. While classical OT has found applications in tasks like domain adaptation [17, 24] and shape matching [22, 40, 77], it often requires solving computationally intensive optimization problems [75, 59]. Schrödinger Bridges [71, 41] extend OT by introducing a stochastic reference process, which enables more flexible and entropically regularized transport. This has proven particularly useful in generative modeling [18, 84, 13, 87, 73, 43, 78], where SBs offer a probabilistic interpolation between data distributions. Recent advances in simulation-free training methods [46, 78, 79] have made these techniques more scalable by bypassing the need to simulate stochastic processes during training. In robotics, these frameworks have been explored for motion planning [39], distribution alignment [76], and dynamical system modeling [9]. However, their application to grasp synthesis and translation remains underexplored. Our work leverages SBs not just for smooth transport between distributions, but for preserving task-specific physical properties of grasps across different hand morphologies, pushing the boundaries of OT/SB utility in robotic manipulation.
@@ 3.1 Schrödinger Bridge Formulation
The Schrödinger Bridge (SB) problem [71, 41] provides a principled framework for defining the most likely stochastic evolution between two probability distributions under a reference process. Given source and target distributions $q_{0}(x)$ and $q_{1}(x)$, and a reference stochastic process $\mathbb{Q}$, the SB seeks a process $\mathbb{P}$ that minimizes the relative entropy (Kullback-Leibler divergence) with respect to $\mathbb{Q}$, while ensuring that $\mathbb{P}$ matches the marginals $q_{0}$ and $q_{1}$ at times 0 and 1, respectively:
Typically, $\mathbb{Q}$ is chosen as a Brownian motion with a diffusion rate $\sigma$, leading to the so-called diffusion Schrödinger Bridge [18, 10]. The optimal process $\mathbb{P}^{*}$ can be characterized as a mixture of Brownian bridges weighted by an entropically regularized optimal transport (OT) plan. Specifically, the entropic OT plan $\pi_{\varepsilon}^{*}(q_{0},q_{1})$ minimizes a regularized cost:
@@
where $U(q_{0},q_{1})$ denotes the set of couplings between $q_{0}$ and $q_{1}$, and $d(\cdot,\cdot)$ is a ground cost. Föllmer [28] connects this formulation to the SB problem: if the reference process $\mathbb{Q}$ is Brownian motion, the SB is uniquely characterized by a mixture of Brownian bridges drawn according to the entropic OT plan. This insight allows the complex dynamic optimization of SB to be reduced to computations over static distributions, making it more tractable in high-dimensional settings.
@@ 3.2 Score-based and Flow-Matching Methods
To model the stochastic dynamics of the SB, we consider Itô SDEs of the form:
@@
where $u_{t}(x)$ is a drift vector field, $g(t)$ is a diffusion schedule, and $w_{t}$ is standard Brownian motion. The marginal densities $p_{t}(x)$ induced by this SDE evolve according to the Fokker-Planck equation [68]:
where $\Delta p_{t}=\nabla\cdot(\nabla p_{t})$ is the Laplacian. An equivalent deterministic process can be described via the probability flow ODE:
@@
which shares the same time marginals $p_{t}$ as the original SDE. The drift term $u_{t}(x)$ can be recovered from the flow drift $u_{t}^{\circ}(x)$ and score $\nabla\log p_{t}(x)$ as $u_{t}(x)=u_{t}^{\circ}(x)+\frac{g(t)^{2}}{2}\nabla\log p_{t}(x)$.
Learning such stochastic dynamics from data typically involves estimating both $u_{t}^{\circ}(x)$ and the score $\nabla\log p_{t}(x)$. In our work, we follow the [SF]2[78] formulation, which proposes to learn these quantities directly by regressing neural networks $v_{\theta}(t,x)$ and $s_{\theta}(t,x)$ to approximate them. In practice, the marginals $p_{t}$ are often not directly accessible. However, by leveraging the connection to entropic OT, one can construct conditional paths (e.g., Brownian bridges) between samples $(x_{0},x_{1})$ from the entropic OT plan and use these to derive analytic expressions for $u_{t}^{\circ}(x|x_{0},x_{1})$ and $\nabla\log p_{t}(x|x_{0},x_{1})$. Specifically, given a source-target pair $(x_{0},x_{1})$ sampled from the entropic OT plan $\pi^{*}_{\varepsilon}$, the conditional distribution at time $t$ is a Gaussian $p_{t}(x\mid x_{0},x_{1})=\mathcal{N}(x;(1-t)x_{0}+tx_{1},\sigma^{2}t(1-t)I)$ with constant diffusion rate $g(t)=\sigma$. From this, we obtain closed-form expressions for the conditional score and probability flow drift:
$u_{t}^{\circ}(x|x_{0},x_{1})=\frac{1-2t}{1-t}(x-((1-t)x_{0}+tx_{1}))+(x_{1}-x_{0}),\ \nabla\log p_{t}(x|x_{0},x_{1})=\frac{(1-t)x_{0}+tx_{1}-x}{\sigma^{2}t(1-t)}.$
@@
$\mathcal{L}_{\mathrm{[SF]}^{2}\mathrm{M}}(\theta)=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}t\sim\mathcal{U}(0,1)\\ x\sim p_{t}(x|x_{0},x_{1})\\ (x_{0},x_{1})\sim\pi^{*}\end{subarray}}\left[\left\lVert v_{\theta}(t,x)-u_{t}^{\circ}(x|x_{0},x_{1})\right\rVert^{2}+\lambda(t)^{2}\left\lVert s_{\theta}(t,x)-\nabla\log p_{t}(x|x_{0},x_{1})\right\rVert^{2}\right],$
where $\lambda(\cdot)$ is some hyperparameter choice of positive weight. We direct readers to the appendix for more details on this formulation.
@@ 4.1 Latent Score and Flow Matching
We approach grasp translation as stochastic transport in a learned latent space where the high-dimensional grasp configurations for source and target hands are first embedded into a shared latent representation. This is achieved through a two-stage pipeline. In the first stage, we train a variational autoencoder (VAE) [38] that encodes segmented visual observations of the source hand grasping an object into a latent variable $z\in\mathcal{Z}$. The VAE consists of an encoder mapping $\mathcal{O}\rightarrow\mathcal{Z}$ and a decoder mapping $\mathcal{Z}\rightarrow\mathcal{G}$. The decoded grasp configuration $\hat{g}=\mathrm{Dec}(z)$ is passed through a differentiable kinematic layer $\mathrm{FK}:\mathcal{G}\rightarrow\mathbb{R}^{3N}$, yielding the 3D mesh vertices of the hand $\hat{v}=\mathrm{FK}(\hat{g})$. The VAE is optimized with a combined loss with a small KL term:
@@
Following Tong et al. [78], we train neural networks $v_{\theta}(t,z)$ and $s_{\theta}(t,z)$ to approximate the conditional flow $u_{t}^{\circ}(z|z_{0},z_{1})$ and score $\nabla\log p_{t}(z|z_{0},z_{1})$, respectively. The training is performed using conditional Brownian bridge samples $z^{t}\sim p_{t}(z|z_{0},z_{1})$ where $(z_{0},z_{1})\sim\pi^{*}_{\varepsilon}$, which is the entropic OT plan defined in Eq. 2 between $q_{0}$ and $q_{1}$. The training objective is then defined as Eq. 7.
@@
1:Trained $\mathrm{Enc}$, source and target dataset, OTPlan, initial $v_{\theta}$,$s_{\theta}$
@@
3:    Sample $(o_{s},g_{s})$, $(o_{t},g_{t})$ from dataset
@@
13:return $v_{\theta}$,$s_{\theta}$
@@
2:$z_{0}\leftarrow\mathrm{Enc}(o_{s})$
3:$z_{1}\sim p_{1}(z|o_{s})$ using $v_{\theta}$, $s_{\theta}$
4:$\hat{g}_{t}\leftarrow\mathrm{Dec}(z_{1})$
@@ 4.2 Grasp-specific OT Cost Design
To enable meaningful grasp translation between different hand morphologies, it is crucial to guide the Schrödinger Bridge process with a ground cost that reflects task-relevant physical properties rather than naive geometric alignment. Conventional entropic OT approaches [78, 79] typically define the ground cost $d(\cdot,\cdot)$ using Euclidean or geodesic distances in either configuration or latent space. However, such metrics fail to capture the functional equivalence of grasps, particularly when source and target hands differ significantly in kinematics and contact modalities. We propose a set of physics-informed cost functions that better encode grasp quality and intent.
Base Pose Cost. We define a Euclidean cost $d_{\mathrm{pose}}$ in $\mathbb{R}^{9}$ between the global poses of the source and target hand. The pose is represented as a concatenation of the 3D base position and a 6D continuous representation of the rotation [94] to ensure smoothness and avoid ambiguities. This cost encourages coarse alignment of the grasps in workspace coordinates:
$d_{\mathrm{pose}}=\left\lVert h_{\mathrm{source}}-h_{\mathrm{target}}\right\rVert^{2}+\left\lVert R_{\mathrm{source}}-R_{\mathrm{target}}\right\rVert^{2},$ (9)
where $h_{\mathrm{source}}$, $h_{\mathrm{target}}$, $R_{\mathrm{source}}$, $R_{\mathrm{target}}$ are translation and rotation of the base pose extracted from $g_{\mathrm{source}}$, $g_{\mathrm{target}}$, respectively.
@@ 4.3 Implementation Details
For the Grasp Wrench Space Overlap cost, we reduce the 6D convex hull to its first three spatial dimensions, approximating the set of achievable center-of-mass forces. Monte Carlo estimation is used to evaluate the IoU of these 3D hulls efficiently on GPUs. For the Jacobian-based maximal effect descriptor, we leverage the Warp differentiable simulator [52] to simulate the grasp and compute the Jacobian of object pose with respect to joint angles. At inference time, latent samples are evolved using Euler–Maruyama integration of the learned SDE with discretized steps. No test-time finetuning is performed. Additional architectural and hyperparameter details are provided in the supplementary materials.

### v2

@@ Abstract
We propose a new approach to vision-based dexterous grasp translation, which aims to transfer grasp intent across robotic hands with differing morphologies. Given a visual observation of a source hand grasping an object, our goal is to synthesize a functionally equivalent grasp for a target hand without requiring paired demonstrations or hand-specific simulations. We frame this problem as a stochastic transport between grasp distributions using the Schrödinger Bridge formalism. Our method learns to map between source and target latent grasp spaces via score and flow matching, conditioned on visual observations. To guide this translation, we introduce physics-informed cost functions that encode alignment in base pose, contact maps, wrench space, and manipulability. Experiments across diverse hand-object pairs demonstrate that our approach generates stable, physically grounded grasps with strong generalization. This work enables semantic grasp transfer for heterogeneous manipulators and bridges vision-based grasping with probabilistic generative modeling. Additional details at grasp2grasp.github.io.
@@ 1 Introduction
Robotic grasping remains a central challenge in autonomous manipulation, especially in the context of dexterous, multi-fingered hands. Unlike parallel-jaw grippers or simple suction devices, dexterous hands offer significantly more control and flexibility but also bring high-dimensional configuration spaces [3, 18] and non-trivial contact dynamics [65]. These properties pose difficulties [53] in data collection, modeling, and learning robust grasp strategies. While recent data-driven works [46, 52, 91, 9, 100, 101] have shown success in generating feasible grasps from large-scale datasets [46, 87, 94, 32], a key limitation persists: models typically operate in a hand-specific setting. This lack of generalization requires retraining or significant fine-tuning to adapt to new robotic hands [41], which poses a practical bottleneck in the rapidly evolving landscape of robotic hardware and limits the applicability of learned policies in real-world systems [97].
Therefore, it is desirable to reuse the grasp knowledge acquired from one hand to infer meaningful grasp strategies for another. However, the morphological differences between hands mean that naive transfer, such as copying joint angles or end-effector poses, might lead to physically invalid or unstable grasps [68, 58, 2]. A more fundamental challenge is the acquisition of paired demonstration data, where a grasp for a source hand is mapped to a functionally equivalent grasp for a target. Manually creating or algorithmically discovering such pairings at scale is extremely difficult. This motivates the need for a framework that can learn to translate the intent of a grasp from a source hand to a target hand using only unpaired, hand-specific grasp datasets, leveraging the underlying object geometry and physics to bridge the morphological gap.
@@ 2 Related Work
Grasp Transfer is closely related to teleoperation [72] and imitation learning [4], where the objective is to map motions or actions from a source agent (e.g., a human or a robot) to a target robot. Teleoperation methods [78, 80, 104, 37, 24, 70] often aim for one-to-one replication of joint trajectories or end-effector poses, which is generally effective when the source and target share similar kinematics. However, such approaches struggle when transferring grasps between manipulators with significantly different morphologies, as direct replication can lead to infeasible or suboptimal grasps [68, 58, 2]. To address this morphological gap, retargeting methods have been explored. For instance, works like [72, 35, 14] explore positional optimization for key points, while CrossDex [98] uses a neural network trained with data generated from [72] for retargeting. These approaches, however, often rely on pre-defined link-to-link correspondences or human-annotated templates, which can limit their generalizability. More recent work, such as RobotFingerPrint [41], addresses this challenge by mapping grippers into a shared Unified Gripper Coordinate Space (UGCS) based on spherical coordinates to facilitate grasp transfer. While promising, this approach requires simulation-based preprocessing and detailed gripper models, which impose a strong prior that limits flexibility. Moreover, the UGCS approach focuses on point correspondences and is thus object-agnostic, rather than explicitly preserving functional grasp intent. In contrast, our formulation aims to translate grasps in a distributional sense grounded in physical properties. It requires no hand-specific simulation preprocessing, which enables broader applicability and more semantically meaningful grasp transfer.
@@
Optimal Transport (OT) and Schrödinger Bridges (SB) have recently emerged as powerful frameworks in the machine learning community for modeling distributional transformations. OT [75, 92, 69, 81] provides a principled way to compute mappings between probability distributions that minimize a cost function, typically grounded in geometric distances. Aligning distributions is a central challenge in tasks like domain adaptation [102, 17], and classical OT has been successfully applied for this purpose [20, 27], as well as for other applications like shape matching [25, 44, 83]. However, it often requires solving computationally intensive optimization problems [81, 63]. Schrödinger Bridges [77, 45] extend OT by introducing a stochastic reference process, which enables more flexible and entropically regularized transport. This has proven particularly useful in generative modeling [21, 90, 15, 93, 79, 47, 84], where SBs offer a probabilistic interpolation between data distributions. Recent advances in simulation-free training methods [50, 84, 85] have made these techniques more scalable by bypassing the need to simulate stochastic processes during training. In robotics, these frameworks have been explored for motion planning [43], distribution alignment [82], and dynamical system modeling [10]. However, their application to grasp synthesis and translation remains underexplored. Our work leverages SBs not just for smooth transport between distributions, but for preserving task-specific physical properties of grasps across different hand morphologies, pushing the boundaries of OT/SB utility in robotic manipulation.
@@ 3.1 Schrödinger Bridge Formulation
The Schrödinger Bridge (SB) problem [77, 45] provides a principled framework for defining the most likely stochastic evolution between two probability distributions, a source $q_{0}(x)$ and a target $q_{1}(x)$, under a reference process. The optimal process can be found by weighting paths according to an entropically regularized optimal transport (OT) plan, $\pi_{\varepsilon}^{*}$. This plan seeks a coupling between the distributions that minimizes a ground cost $d(\cdot,\cdot)$ plus a regularization term:
@@
where $U(q_{0},q_{1})$ is the set of all couplings between $q_{0}$ and $q_{1}$. This connection allows the complex dynamic optimization to be grounded in a more tractable static transport plan. We refer readers to Appendix A for a detailed formulation.
@@ 3.2 Score-based and Flow-Matching Methods
We model the stochastic dynamics of the SB with an Itô SDE of the form:
@@
where $u_{t}(x)$ is a drift vector field and $w_{t}$ is standard Brownian motion. This process induces time marginals $p_{t}(x)$ that can also be described by an equivalent deterministic probability flow ODE:
@@
Following the [SF]2M framework [84], we learn the stochastic dynamics by training neural networks $v_{\theta}(t,x)$ and $s_{\theta}(t,x)$ to approximate the ODE drift ($u_{t}^{\circ}$) and the score term ($\nabla\log p_{t}(x)$), respectively. The training targets are derived from conditional Brownian bridge paths between pairs of samples $(x_{0},x_{1})$ drawn from the entropic OT plan $\pi_{\varepsilon}^{*}$ (Eq. 1). A visualization of these conditional targets is provided in Fig. 2.
Figure 2: Illustration of the Schrödinger Bridge Models. The Schrödinger Bridge process transports samples from a source distribution ($q_{0}$) to a target ($q_{1}$). At an intermediate point $x_{t}$ on a Brownian bridge trajectory between coupled samples $(x_{0},x_{1})$, our model learns the conditional flow drift $u_{t}^{\circ}$ that drives transport and the conditional score $\nabla\log p_{t}$ that corrects the path.
@@
$\mathcal{L}_{\mathrm{[SF]}^{2}\mathrm{M}}(\theta)=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}t\sim\mathcal{U}(0,1)\\ x\sim p_{t}(x|x_{0},x_{1})\\ (x_{0},x_{1})\sim\pi_{\varepsilon}^{*}\end{subarray}}\left[\left\lVert v_{\theta}(t,x)-u_{t}^{\circ}(x|x_{0},x_{1})\right\rVert^{2}+\lambda(t)^{2}\left\lVert s_{\theta}(t,x)-\nabla\log p_{t}(x|x_{0},x_{1})\right\rVert^{2}\right],$
where $u_{t}^{\circ}(\cdot)$ and $\nabla\log p_{t}(\cdot)$ are the conditional ODE drift and score targets, and $\lambda(t)$ is a weighting schedule. The full derivation of the training targets and weighting schedule is provided in Appendix A.
@@ 4.1 Latent Score and Flow Matching
[…]
@@
As introduced in Sec. 3.2, we train neural networks $v_{\theta}(t,z)$ and $s_{\theta}(t,z)$ to approximate the conditional flow $u_{t}^{\circ}(z|z_{0},z_{1})$ and score $\nabla\log p_{t}(z|z_{0},z_{1})$, respectively. The training is performed using conditional Brownian bridge samples $z^{t}\sim p_{t}(z|z_{0},z_{1})$ where $(z_{0},z_{1})\sim\pi^{*}_{\varepsilon}$, which is the entropic OT plan defined in Eq. 1 between $q_{0}$ and $q_{1}$. The training objective is then defined as Eq. 4.
@@
1:Trained $\mathrm{Enc}$, source and target dataset, OTPlan, initial $v_{\theta}$$s_{\theta}$
@@
3:  Sample $o_{s},o_{t}$ from dataset
@@
13:return $v_{\theta}$$s_{\theta}$
@@
[…]
@@ 4.2 Grasp-specific OT Cost Design
To guide the Schrödinger Bridge, we must first compute an entropic OT plan $\pi_{\varepsilon}^{*}$ (Eq. 1), which is defined by a ground cost $d(\cdot,\cdot)$. We design several physics-informed costs for this purpose. For a given cost, we compute the OT plan between minibatches of source and target latent samples. We then draw pairs $(z_{0},z_{1})\sim\pi_{\varepsilon}^{*}$ and use these pairs to construct the conditional training targets for the SB model as described in Sec. 3.2 and used in the loss function of Eq. 4. This allows us to enable meaningful grasp translation between different hand morphologies by guiding the process with a ground cost that reflects task-relevant physical properties rather than naive geometric alignment. Conventional entropic OT approaches [84, 85] typically define the ground cost $d(\cdot,\cdot)$ using Euclidean or geodesic distances in either configuration or latent space. However, such metrics fail to capture the functional equivalence of grasps, particularly when source and target hands differ significantly in kinematics and contact modalities. We propose a set of physics-informed cost functions that better encode grasp quality and intent.
Base Pose Cost. We define a cost $d_{\mathrm{pose}}$ on the global wrist pose in $SE(3)$ between the source and target hands. The pose is represented as a concatenation of the 3D base position and a 6D continuous representation for the rotation [103] to ensure smoothness. The 6D vector is first converted to a $3\times 3$ rotation matrix $R\in SO(3)$. The cost then combines the squared L2 norm for translation and the squared Frobenius norm for rotation:
$d_{\mathrm{pose}}=\left\lVert h_{\mathrm{source}}-h_{\mathrm{target}}\right\rVert_{2}^{2}+\left\lVert R_{\mathrm{source}}-R_{\mathrm{target}}\right\rVert_{F}^{2},$ (6)
where $h\in\mathbb{R}^{3}$ is the translation vector and $R\in SO(3)$ is the rotation matrix of the base pose, extracted from $g_{\mathrm{source}}$ and $g_{\mathrm{target}}$, respectively. This cost encourages coarse alignment of the grasps in workspace coordinates.
@@ 4.3 Implementation Details
For the Grasp Wrench Space Overlap cost, we reduce the 6D convex hull to its first three spatial dimensions, approximating the set of achievable center-of-mass forces. Monte Carlo estimation is used to evaluate the IoU of these 3D hulls efficiently on GPUs. For the Jacobian-based maximal effect descriptor, we leverage the Warp differentiable simulator [56] to simulate the grasp and compute the Jacobian of object pose with respect to joint angles. At inference time, latent samples are evolved using Euler-Maruyama integration of the learned SDE with discretized steps. No test-time finetuning is performed. Additional architectural and hyperparameter details are provided in the supplementary materials.